presmic:

The hardcover version of a book weighs 7 ounces while its paperback version weighs 5 ounces. Forty-five copies of the book weigh a total of 249 ounces. A table titled Weight of Books showing Number of Books, Weight per book in ounces, and Total Weight in ounces. The first row shows Hard Cover, with h, 7, and 7 h. The second row shows Paperback, with 45 minus h, 5, and x. The third row shows Total, with 45, blank, and 249. Which value could replace x in the table? 40 – h 50 – h 5(45) – h 5(45 – h)

umm:

We'll need to construct a system of two equations with two variables. So, let "h" be the number of hardcover version books, and let "p" be the number of paperback version books. Given the hardcover version of a book weighs 7 ounces, and the paperback version weighs 5 ounces, to reach a total of 249 ounces we should have \[7h+5p=249\] and there are forty-five (45) copies of the book \[h+p=45\] from that, we clear the number of books paperback versions we have\[p=45-h\] and since each paperback version book weighs 5 ounces, to gather the total weight of the paperback version books, represented by "x" with the table shown, we multiply \[5 \times p=5(45-h)\] which brings us down to answer choice D, \[x = 5(45-h)\] @Vocaloid - Please check when able to...